The income from an 8% investment is $300 more than the income from a 6% investment. How much invested at each rate? If a total of $30000 invested?
Answer by dkppathak(439)

The income from an 8% investment is $300 more than the income from a 6% investment. How much invested at each rate? If a total of $30000 invested?
let $ X invested at 8%
and $ 30000-X at 6%
0.08X-.06(30000-X)=300
.08x+0.06x-1800=300
.14x=300+1800
.14x=2100
x=2100/.14=15000
other amount will be 30000-15000=15000
both investment are $15000 at 8%
and investment at 6% also $15000
Answer $15000 at 8%
$15000 at 6%